# MeshUp: Multi-Target Mesh Deformation via Blended Score Distillation

### v2

@@ Appendix A Additional Experiments
In Figure 16, we show that any keyframe concepts generated in our system can be further interpolated easily, thanks to dense correspondences maintained by our mesh deformation technique. Although our method deforms the source mesh into a blend of various shapes on discreet weight scales $w$, we can use our method to define key-point meshes and interpolate these key points to get an even more fine-grained blend of targets.
3D ROI Ablation. We show an ablation of the 3D ROI map $V_{R}$ for our localized deformation method in Figure 17. Due to the viewpoint consistency of our 3D ROI map, our method can generate smooth, meaningful mixing results that respect the specified local regions for each target. This is in stark contrast to the results we get without using the 3D ROI map (directly using the self-attention maps extracted from each rendering to mask cross-attention maps), where we observe sharp artifacts as well as significant loss of details for the specified targets.
Modified Classifier Free Guidance Ablation. In Figure 18, we show an ablation of our modified Classifier Free Guidance for self-blending experiments. We notice that the effect of this modification is particularly crucial on smaller weights (typically from 0% to 15% of target weight) since for the self-blending application, a low target weight implies that there is more bias introduced from the null text prompt. Notice that the results conditioned on smaller weights without our modified classifier free guidance are severely irregular and biased, while using the modified classifier free guidance stabilizes this bias, regulating the deformation in a much stable, smooth manner.
@@
Appendix B Comparisons
Since our method is the first one to address the concept mixing problem in mesh deformation, we create several baselines using existing methods and compare our results to what these baselines can achieve.
Shape Interpolation Baseline. We also consider a simple interpolation baseline. We use our single-branch SDS with Jacobian optimization to create two deformations of the source shape with respect to two different textual targets. Then, we directly interpolate the vertex positions between these two targets, as presented in Figure 19. The shape interpolation baseline does not enable new features to emerge for the in-between shapes, thus, leading to non-descriptive, overly-smooth interpolation results. In contrast, our BSD method clearly prioritizes the emergence of notable features of each target (notice how the legs of the Kangaroo emerge first while the face of the Dachshund is clearly prioritized, respective of the weights given to each target).
We provide a more extensive qualitative comparison for the single-target deformation between TextDeformer and our method in the supplements. Although the main focus of the paper is concept mixing, the figure shows that our method substantially improves the single-target deformation quality over TextDeformer.
@@ Appendix B Comparisons
Table 3: Quantitative evaluation. We compare MeshUp to TextDeformer [18] and report CLIP R-Precision [45]. Note that TextDeformer has an advantage to our method since it is directly supervised with CLIP loss.
We highlight the significant accuracy of the users’ guesses, suggesting that the BSD weights control the concept blending in an intuitive and plausible manner. The set of renderings that we used for both evaluations can be found in the supplementary section.
@@
Quantitative Comparison We compare our method to TextDeformer by running single-target deformation for both methods. We use a source dog mesh and warp it to the following 10 different prompts: “bear”, “bulldog”, “dachshund”, “desertfox”, “frog”, “hippo”, “kangaroo”, “pig”, “puma”, “siberian cat.” We evaluate our result using CLIP R-Precision score and show results in Table  3. Our method outperforms TextDeformer on most metrics. The results that we used for both this evaluation and user study can be found in figure 20.

### v3

@@ Appendix B Additional Experiments
3D ROI Ablation. We show an ablation of the 3D ROI map $V_{R}$ for our localized deformation method in Figure 17. Due to the viewpoint consistency of our 3D ROI map, our method can generate smooth, meaningful mixing results that respect the specified local regions for each target. In contrast, the results we get without using the 3D ROI map (by directly using the 2D ROI map extracted from self-attention maps), where we observe sharp artifacts as well as significant loss of details for the specified targets.
Modified Classifier Free Guidance Ablation. In Figure 18, we show an ablation of our modified Classifier Free Guidance for self-blending experiments. We notice that the effect of this modification is particularly crucial on smaller weights (typically from 0% to 15% of target weight) since for the self-blending application, a low target weight implies more bias from the null text prompt. Notice that the results conditioned on smaller weights without our modified classifier free guidance are severely irregular and biased, while using the modified classifier free guidance stabilizes this bias, regulating the deformation in a much stable, smooth manner.
@@
Appendix C Comparisons
Since our method is the first to address the concept mixing in mesh deformation, we create simple baselines using existing methods and compare our results to what these baselines can achieve in both qualitative and quantitative metrics.
Shape Interpolation Baseline. We consider a simple interpolation baseline, where we generate deformations of two distinct targets using our single target SDS baseline, and directly interpolate the vertex positions between the two meshes. We then compare this naive vertex-wise interpolation baseline to results generated from our BSD method in Figure 19 We notice that the shape interpolation baseline does not enable new features to emerge for the interpolated shapes, while our method clearly prioritizes the emergence of notable features of each target (notice how the legs of the Kangaroo emerge first while the face of the Dachshund is clearly prioritized, respective of the weights given to each target).
@@ Appendix C Comparisons
Table 3: Quantitative evaluation. We compare MeshUp to TextDeformer [18] and report CLIP R-Precision [46] scores.
We highlight the significant accuracy of the users’ guesses, suggesting that the BSD weights control the concept blending in an intuitively plausible manner. The set of renderings that we use for both evaluations can be found in Figure 20.
@@
Quantitative Comparison We use the same dataset to quantitatively compare our method to TextDeformer using CLIP R-Precision Scores. We use a source dog mesh and warp it to the following 10 different prompts: “bear”, “bulldog”, “dachshund”, “desertfox”, “frog”, “hippo”, “kangaroo”, “pig”, “puma”, “siberian cat.” We evaluate our result using CLIP R-Precision score and show results in Table  3. Our method outperforms TextDeformer on most metrics, despite the metric being inherently favorable to TextDeformer, since its optimization is directly supervised by CLIP score.